OpenStudy (anonymous):

. If Lynn can type a page in p minutes, what piece of the page can she do in 5 minutes? What would the answer be ? A. 5/p B. p - 5 C. p + 5 D. p/5 E. 1- p + 5 I really don't understand this can someone please help me

myininaya (myininaya):

1 page =p minute 1/p pages=1 minute 5(1/p pages)= 5(1 minute) 5/p pages=5 minutes

myininaya (myininaya):

make sense?

OpenStudy (anonymous):

Yesss , way better now thank you !